-7(y-12)=-147 what is y?
Found 2 solutions by erica65404, Alan3354:
Answer by erica65404(394)
Answer by Alan3354(69443)

-7(y-12)=-147 what is y?
-----------
y is a variable.
---------
If you want to solve for y, you should say, "Solve for y."
-7(y-12)=-147
7(y-12) = 147
Divide by 7
y-12 = 21
Add 12
y = 33